f(x)=-2x^2-3x+8, find f(-2)please!
Answer by mifarnc(2)

If the f(x) is equal to -2 Then,
replace -2 for x in the problem.
Such as:
f(x)= -2x^2-3x+8
F(-2)= -2(-2)^2-3(-2)+8
rememebring your order of operations, complete the exponent first.
(-2)^2 = 4
-2(4)-3(-2)+8
Use order of operations again
-2x-4=-8
-3x-2=6
so now we should have
-8+6+8
then solve
Therefore the F(-2) is 6